Carreras Café is a Spanish restaurant in a college town. The owner expects that the number of meals served in June will be 40 percent below those served in May, because so many students leave for the summer. In May, the restaurant served 4,200 meals at an average price of $20. In the summer (June through August), the average price of a meal typically increases by 30 percent and the average food cost of a meal increases by 15 percent. Other costs are typically unchanged during the summer. The following cost information is available for May.


Service costs (variable with respect to meals)  

Variable service costs (with respect to meals)  

Food costs                                                       $10,200

Labor                                                                 16,800

Other variable costs                                             5,900

Fixed service costs                                               9,300

Total service costs                                              $42,200

Marketing and administrative costs  

Marketing (variable with respect to meals)          6,400

Administrative (fixed)                                            4,900

Total marketing and administrative costs           $11,300

Total costs                                                          $53,500


Required:

Prepare a budgeted income statement for June. (Do not round intermediate calculations.)
Answer #1

1-.04= 0.6            1+.30= 1.30    1+0.15= 1.15

Sales revenue (2,520 meals @ $26.00/meal)   ($20 × 4,200 × 0.60 × 1.30)   $65,520

Food cost ($10,200 × 0.60 × 1.15)= 7,038

Labor  ($16,800 × 0.60)= 10,080

Other variable costs ($5,900 × 0.60)= 3,540

Fixed service costs (unchanged)= 9,300

Marketing (variable)=($6,400 × 0.60)= 3,840

Administrative (fixed) (unchanged) 4,900


20210325_133911.jpg